of what the answer to.
the sides of a triangle have these lengths :2x+3,4x+1x+7 the perimeter is 25in. find the length of each side using the formula p=a+b+c

Answer by ewatrrr(24785)


Hi,
Am asssuming the lengths of the sides are represented by:
2x+3,4x+1 and x+7
Perimeter(sum of the lengths of the sides) is 25in. find the length of each side
2x+3 + 4x+1 + x+7 = 25in
solving for x
7x + 11 = 25in
7x = 14in
x = 2in
Length of the sides are:
4+3 = 7in
8+1 = 9in
2 + 7 = 9in
checking our answer
7in + 9in + 9in = 25in